# OpenCodeReasoning: Advancing Data Distillation for Competitive Coding

### v1

@@ 3 Main Evaluation
We assessed the effectiveness of SFT using OpenCodeReasoning by fine-tuning the Qwen2.5 base and instruct models at 7B, 14B, and 32B parameters. The models were trained for 3 epochs on NVIDIA H100-80GB GPUs, using the AdamW optimizer (Kingma & Ba, 2015) with a batch size of 256 and a maximum sequence length of 32,768. An initial learning rate grid search, covering $[1e-5,3e-5,5e-5,8e-5,1e-4]$, identified $55e-$ 5 as optimal across model sizes. We employed a CosineAnnealing scheduler with a warmup ratio of 0.1, and used the final checkpoint for evaluation. Training acceleration was achieved through sequence packing (Shen et al., 2024), tensor and context parallelism, and BF16 precision. For inference, we employed temperature-based nucleus sampling (Holtzman et al., 2020). An initial temperature sweep over the values $[0.0,0.2,0.6,0.7,1.0]$ revealed that temperatures of 0.6 and 0.7 yielded the best performance. Consequently, we selected a temperature of 0.6 for all subsequent experiments. Inference was performed using vLLM (Kwon et al., 2023) with a maximum generation length of 30,720 tokens.
@@ 4.2 Ablation: Inclusion of C++ Solutions
While OpenCodeReasoning primarily consists of Python samples, we sought to investigate whether incorporating solutions in other languages could enhance distillation. Consequently, we explored the inclusion of 356k C++ samples, generated by R1, in conjunction with OpenCodeReasoning. In addition to LiveCodeBench and CodeContest, we evaluated on the IOI benchmark (Penedo et al., 2025b), which requires C++ solutions. The results are presented in Table 4. It is evident that the inclusion of C++ solutions has no positive impact on Python benchmark performance, but does significantly improve the accuracy on the C++ benchmark. Finding an optimal strategy to effectively leverage multilingual data warrants further research.
@@ Appendix A Reasoning pattern extraction
Final chain of thought segmentation prompt Below is a chain of thought for solving a question. For the segment between the $<$unannotated$>$ and </absent</< /unannotated$>$ tags, figure out what is reasoning pattern used in that segment like problem rephrasing, new idea generation, self-evaluation, verification, backtracking, subgoal generation, solving a simpler problem first, or something else. Then your task is to identify the reasoning pattern used in the unannotated segment. Rewrite the unannotated segment in the following format:
@@
We use Qwen-32B-Instruct model to do the initial segmentation of the chain of thought using the prompt showed in Figure 6. Then for each unannotated segment, we use the prompt in Figure 7 to do one more round of segmentation. The reasoning patterns are then extracted from within the $<$pattern$>$ tags. We merge the verification and reasoning pattern into self-evaluation itself. Anytime the model labels a segment with multiple patterns, we exclude that segment as we consider the model unsure of what pattern it corresponds to. For each generation we compute the fraction of times a particular pattern appears in the generation out of all the patterns in the generation. Thus for each generation we obtain a vector with element corresponding to how frequently that pattern appeared in the generation. For the analysis of patterns with respect to hardness, we computed the mean of this distribution across all correct and incorrect generations separately, using a t-test to compute significance. To analyze per-problem pattern prevalence, we generated a binary matrix where each row represented a problem and each element indicated whether a specific pattern was more frequent in correct solutions (1) or incorrect solutions (0). Statistical significance was then determined using the binomial test.

### v2

@@ 3 Main Evaluation
We assessed the effectiveness of SFT using OpenCodeReasoning by fine-tuning the Qwen2.5 base and instruct models at 7B, 14B, and 32B parameters. The models were trained for 3 epochs on NVIDIA H100-80GB GPUs, using the AdamW optimizer (Kingma & Ba, 2015) with a batch size of 256 and a maximum sequence length of 32,768. An initial learning rate grid search, covering [1e5,3e5,5e5,8e5,1e4][1e-5,3e-5,5e-5,8e-5,1e-4], identified 5e55e-5 as optimal across model sizes. We employed a CosineAnnealing scheduler with a warmup ratio of 0.1, and used the final checkpoint for evaluation. Training acceleration was achieved through sequence packing (Shen et al., 2024), tensor and context parallelism, and BF16 precision. For inference, we employed temperature-based nucleus sampling (Holtzman et al., 2020). An initial temperature sweep over the values $[0.0,0.2,0.6,0.7,1.0]$ revealed that temperatures of 0.6 and 0.7 yielded the best performance. Consequently, we selected a temperature of 0.6 for all subsequent experiments. Inference was performed using vLLM (Kwon et al., 2023) with a maximum generation length of 30,720 tokens.
@@ 4.2 Ablation: Inclusion of C++ Solutions
While OpenCodeReasoning primarily consists of Python samples, we sought to investigate whether incorporating solutions in other languages could enhance distillation. Consequently, we explored the inclusion of 356k C++ samples, generated by R1, in conjunction with OpenCodeReasoning. In addition to LiveCodeBench and CodeContest, we evaluated on the IOI benchmark (Penedo et al., 2025b), which requires C++ solutions. The results are presented in subsection 4.2. It is evident that the inclusion of C++ solutions has no positive impact on Python benchmark performance, but does significantly improve the accuracy on the C++ benchmark. Finding an optimal strategy to effectively leverage multilingual data warrants further research.
@@ Appendix A Reasoning pattern extraction
Final chain of thought segmentation prompt Below is a chain of thought for solving a question. For the segment between the <<unannotated>> and </</unannotated>> tags, figure out what is reasoning pattern used in that segment like problem rephrasing, new idea generation, self-evaluation, verification, backtracking, subgoal generation, solving a simpler problem first, or something else. Then your task is to identify the reasoning pattern used in the unannotated segment. Rewrite the unannotated segment in the following format:
@@
We use Qwen-32B-Instruct model to do the initial segmentation of the chain of thought using the prompt showed in Figure 6. Then for each unannotated segment, we use the prompt in Figure 7 to do one more round of segmentation. The reasoning patterns are then extracted from within the <<pattern>> tags. We merge the verification and reasoning pattern into self-evaluation itself. Anytime the model labels a segment with multiple patterns, we exclude that segment as we consider the model unsure of what pattern it corresponds to. For each generation we compute the fraction of times a particular pattern appears in the generation out of all the patterns in the generation. Thus for each generation we obtain a vector with element corresponding to how frequently that pattern appeared in the generation. For the analysis of patterns with respect to hardness, we computed the mean of this distribution across all correct and incorrect generations separately, using a t-test to compute significance. To analyze per-problem pattern prevalence, we generated a binary matrix where each row represented a problem and each element indicated whether a specific pattern was more frequent in correct solutions (1) or incorrect solutions (0). Statistical significance was then determined using the binomial test.